OpenStudy (anonymous):

Large truck with massof 1600kg pulling a trailer with mass of 1200kg. Both accelerate at 0.70m/s^2. Determine the tension force between the truck and trailer and the force with which the truck is pulling if the force of friction is 250 N for the trailer and 500 N for the truck. I dont get how to solve for tension at all, I really need help. Please tell me how to solve for tension simply?

OpenStudy (anonymous):

Given the forces of friction, the masses and acceleration, you can use F = ma.

OpenStudy (anonymous):

your a fluttering dumbass -.-

OpenStudy (anonymous):

F - T - 500 = 1600 * 0.70 F + T -250 = 1200 * 0.70  Work out what T and F are.